OpenStudy (anonymous):

solve the equation 3tan^2x-1=0 for 0<_ x <_360degree. give your answer to the nearest degree

jhonyy9 (jhonyy9):

3tan*2x-1=0 us tan*2x can be writing like (sin*2 x)/(cos*2 x) hence you get (3sin*2 x)/cos*2 x -1=0 --- 3sin*2 x -cos*2 x =0 but you know that sin*2 x + cos*2 x =1 from what cos*2 x=1-sin*2 x that your equation will be 3sin*2 x -(1-sin*2 x)=0 3sin*2 x -1+sin*2 x=0 --- 4sin*2 x - 1 = 0 if you sign sin x=a you will get 4a*2 -1 =0 here you can using a*2 - b*2 =(a+b)(a-b) than 4a*2 -1 =(2a-1)(2a+1)=0 --- 2a-1=0 -- 2a=1 -- a=1/2 2a+1=0 -- 2a=-1 --- a=-1/2